OpenStudy (ledah):

Simplify the following (-8a^3b^4)2

OpenStudy (unklerhaukus):

\[(-8a^3b^4)^2=(-8)^2a^{3\times2}b^{4\times2}=\cdots\]

OpenStudy (ledah):

how do I simplify if there are no other terms besides the paren- nevermind o_o

OpenStudy (unklerhaukus):

remember negative times negative is positive

OpenStudy (ledah):

I dont think this is right but I got -8a^6b^8

OpenStudy (ledah):

oh wait positive 8... sorry accidentally hit the "-" sign

OpenStudy (unklerhaukus):

almost the only bit you got wrong was \[(-8)^2=-8\times-8\neq-8\]

OpenStudy (unklerhaukus):

\[(-8)^2=-8\times-8\neq8\]

OpenStudy (ledah):

8x8? wouldn't that be 64?

OpenStudy (unklerhaukus):

\[(-8)^2=-8\times-8=8\times8=\cdots\]

OpenStudy (unklerhaukus):

yeap, you got it

OpenStudy (ledah):

thanks.

OpenStudy (unklerhaukus):

what was your final answer